Imelda got a score of 98,16,17,89,22,14,21,34,56 in her last quiz in midterm. What is her average score?

1
Expert's answer
2021-11-08T20:31:35-0500

"98,16,17,89,22,14,21,34,56"

"\bar{x}=\dfrac{1}{n}\displaystyle\sum_{i=1}^nx_i"

"\bar{x}=\dfrac{1}{9}(98+16+17+89+22+14"

"+21+34+56)=\dfrac{367}{9}\approx40.8"

Her average score is 40.8.